OpenStudy (anonymous):

Witch fraction is equivilent to 5 3/4

OpenStudy (gorv):

5+3/4

OpenStudy (anonymous):

witch is?

OpenStudy (anonymous):

5 3/4 = 23/4

OpenStudy (gorv):

@RyleeC is right

OpenStudy (anonymous):

anser more for me ryleeC

OpenStudy (anonymous):

Lets see.

OpenStudy (anonymous):

witch mixed number is eqivilent to 708/100

OpenStudy (anonymous):

7 2/25

OpenStudy (anonymous):

thanks 2 more please

OpenStudy (anonymous):

what is the least common denominater for the fractions 3/16 and 11/24? @RyleeC

OpenStudy (anonymous):

First find the CD

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

3/16 = 72/384 11/24 = 176/384

OpenStudy (anonymous):

so the anser is ?

OpenStudy (anonymous):

now simplify each fraction until they have a equal denominator

OpenStudy (anonymous):

how

OpenStudy (anonymous):

do it for me please

OpenStudy (anonymous):

find common denominator for example 4/8 = 1/2 = 6/12

OpenStudy (anonymous):

I cant do everything.

OpenStudy (anonymous):

please

OpenStudy (anonymous):

im confused @RyleeC

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

what is the least common denominater for the fractions 3/16 and 11/24?

OpenStudy (anonymous):

3/16 and 11/24 cannot be broken down anymore, so we must raise them.

OpenStudy (anonymous):

the goal is to find where their denominators are the same.

OpenStudy (anonymous):

What the answer choices?

OpenStudy (anonymous):

For (16, 24) the least common multiple (LCM) is 48. Therefore, the least common denominator (LCD) is 48. Rewriting as equivalent fractions with the LCD 9/48, 22/48

OpenStudy (anonymous):

these are the ansers A. 24 B. 40 C. 48 D. 384

OpenStudy (anonymous):

so its c.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you can you anser 1 more?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Cynthia made the conjecture that the sum of any prime number and any composite number is a composite number. Which equation is a counterexample to her conjecture?

OpenStudy (anonymous):

A. 11 + 2 = 13 B. 13 + 2 = 15 C. 5 + 8 = 13 D. 11 + 9 = 20

OpenStudy (anonymous):

this case : prime + composite = prime

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

so whats the anser

OpenStudy (anonymous):

abcord

OpenStudy (anonymous):

abc or d

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np